### v1

# Dominique Levesque Laboratoire des 2 infinis Irène Joliot-Curie, Université Paris-Saclay    Nicolas Sourlas Laboratoire de Physique de l’ Ecole Normale Supérieure, ENS, Université PSL, F-75005 Paris, France

Abstract
One of the important questions in statistical mechanics is how irreversibility (time’s arrow) occurs when Newton equations of motion are time reversal invariant. Boltzmann’s argument is that the observed irreversibility is due to the very large number of particles. One objection to Boltzmann was based on Poincaré’s recursion theorem: a classical hamiltonian confined system returns after some time, called today Poincaré recurrence time (PRT), close to its initial configuration. Boltzmann’s reply was that for $N=10^{23}$ particles, PRT is very large and exceeds the age of the universe. In this paper we compute for the first time, a typical recurrence time $T(N)$ for a gas of $N$ particles. We find that $T(N)\sim N^{z}\exp(yN)$ and determine the exponents $y$ and $z$ for different values of the particle density and kinetic energy (temperature). We find that $T(N)$ exceeds the age of the Universe for a relatively small number of particles, much smaller than $10^{23}$. Our results support Boltzmann’s argument on the origin of irreversibility in statistical mechanics.
One of the important questions in statistical mechanics is how time irreversibility occurs (times arrow), despite the fact that the equations of motion obey time reversal invariance. This is a long and until recently debated issue. The typical example in the litterature is the following. You drop a small quantity of coffe in the corner of a glass of milk. Coffe diffuses slowly and the liquid becomes homogenious and stays homogenious (the state of maximal entropy) for as long as you can wait. But according to Poincaré’s recursion theorem[1] a classical hamiltonian confined system, evolving according to the classical equations of motion, returns close to its initial condition after some time $t$. This time $t$ is called today the Poincaré recurrence time (PRT).
Boltzmann’s answer to his apparent paradox is that time irreversibility in macroscopic physical systems is due to their very large number of particles. Starting from any initial condition, a macroscopic system will move toward states of increasing entropy and thermal equilibrium. Boltzmann estimated the PRT for a gas and argued that for $10^{23}$ molecules this time is larger than the life of the universe. For an excellent discussion on the subject see the article of Lebowitz [2] and the book of Gallavotti [3]. The majority but not all of physicists today agree with Boltzmann.
Some people argue that contrary to Boltzmann’s argument, irreversibility is due to elementary interactions which violate time reversal symmetry[5]
To the best of our knowledge the PRT has never being precisely estimated up to now, except for the case of some particular integrable systems [4].
In the present paper we use molecular dynamics to numerically measure $T(N)$, a typical PRT for a system of $N$ Lennard Jones particles. We measure $T(N)$ for a relatively small number of particles $N$ and use finite size scaling arguments to extrapolate towards the limit of large $N$. We find that $T(N)\sim N^{z}\exp(yN)$ and compute the exponents $y$ and $z$ for different values of the particle density and the kinetic energy (temperature) of the gas. More specifically we consider $N$ particles confined in a sphere by a surface potential: $v_{s}(R,r)=A/(|\overrightarrow{R}-\overrightarrow{r}|)^{2n}$ where $\overrightarrow{R}$ is the position of a sphere point and $\overrightarrow{r}$ a particle position. At time $t=0$ all particles are in a half sphere, i.e. their coordinates $x$ are $<0$, and their velocity obeys a gaussian distribution with average $v$. The particles interact via a Lennard Jones potential
$v(r_{ij})=4\epsilon((\frac{\sigma}{r_{ij}})^{12}-(\frac{\sigma}{r_{ij}})^{6})+\epsilon$
if $r_{ij}<2^{1/6}$ and $v(r_{ij})=0$ if $r_{ij}>2^{1/6}$, $\sigma$ and $\epsilon$ being respectively the length and energy units. The standard algorithm for simulating Lennard Jones particles is the Verlet algorithm. We integrate the Newton equations of motion by using a variant of the Verlet algorithm described in reference [6], an algorithm which, based on integer arithmetic, strictly respects the time reversal symmetry. Indeed, it has been shown that rounding-off errors in floating-point arithmetic can induce violations of the time reversal symmetry. A strict respect of time reversal symmetry allows to exlude the argument that irreversibility is due to interactions violating time reversal symmetry [5].
Figure 1: Typical fit of $\log(T(N))$ at density $d$=0.064.
Even for small $N$ systems it would take a too long time in order that the $N$ particles come back almost exactly to their velocities and positions in the initial configuration. Instead we consider $n_{i}(N)$ systems differing by the initial conditions and define for each system a recurrence time $T_{i}(N)$ which is chosen as the time it takes, for the i’th system, so that $43N/$ 4 particles (or more) come back to the same half sphere as when $t=0$, i.e the time separating the occurence of an important density fluctuation in a half sphere which, usually in equilibrium states, contains $\sim N/2$ particles.
Increasing $N$, $v$ being constant, we increase the radius of the sphere to keep the density of particles $d$ constant. We observe that $T_{i}(N)$ strongly fluctuates depending on the initial conditions. For every $N$, we average $T_{i}(N)$ over a large number $n_{i}(N)$ of initial conditions and compute $T(N)$, the $T_{i}(N)$ average over those initial conditions. For $N<25$ we simulate $n_{i}(N)=10000$ samples, while for the largest number of $N$ we simulate, which depends on the density $d$, $n_{i}(N)\sim 1000$. We found that the dependence of $T(N)$ on $N$ is exponential and is well described by the formula $T(N)=c^{{}^{\prime}}N^{z}\exp(yN)$ or $\log(T(N))=yN+z\log(N)+c$ as it is shown in Fig. 1.
Figure 2: Region of $y$ and $z$ exponents with 90% confidence level, $d$=0.064 and $v$=2.0
We compute the exponents $y$ and $z$. $y$ and $z$ are correlated. This is illustrated in Fig. 2 which shows their allowed values with 90 percent confidence. $y$ and $z$ depend on the density $d$ and the average velocity $v$. The results of these computations are shown in the Table.
In the first column is listed the particle density $d$, in the second the range of the exponents $y$ and $z$ with 90 percent confidence and average velocity $v=2.0$ (cf. Fig. 2), in the third the same for $v=5.0$, in the last column $N_{m}$ is the maximum number of particles we simulated for that density. The quality of the fit is excellent, as shown in Fig 1. with $\chi^{2}$ per degree of freedom $\chi^{2}<1.$ We found that $T(N)$ is exponential in the number of particles. The exponent $y$ increases with the particle density $d$. The dependance on $v$ (the temperature) is weaker.
$\ d$ $v=2.0$ $v=5.0$ $N_{m}$
$0.064$ $0.1780.170<y<$ 0.178 $0.1730.163<y<$ 0.173 $64$
$0.064$ $-0.526<z<-0.286$ $-0.769<z<-0.437$ $64$
$0.127$ $0.2440.218<y<$ 0.244 $0.2270.201<y<$ 0.227 $48$
$0.127$ $-0.916<z<-0.221$ $-1.112<z<-0.447$ $48$
$0.191$ $0.3120.281<y<$ 0.312 $0.2550.233<y<$ 0.255 $48$
$0.191$ $-0.938<z<-0.175$ $-0.748<z<-0.242$ $48$
Table 1: $y$ and $z$ values obtained from fits of $\ln(T(N))$.
Clearly,our definition is diferent from the PRT standard definition which requires recurrence in the phase space, i.e. both for the particle position and momentum. $T(N)$ is only the recurrence time of an important density fluctuation. Nevertheless the algorithm conserves energy and we have verified that the particle velocity distribution stays gaussian, while individual positions and velocities of the particles may differ from their initial value. Hence $T(N)$ corresponds to the recurrence of configurations much less precisely characterized than for PRT, so $T(N)$ should be certainly an underestimation of PRT.
Simulation of systems with larger particle numbers and densities than mentioned in the Table, would take too much computer time, in particular since the exponent $y$ increases with the density.
It is clear that because of its exponential increase with the number of particles, $T(N)$ will become larger than the life of the universe provided that the number of particles is large enough, say for $N>N_{u}$. The current estimate of the life of the universe is  $\sim\,14\times 10^{9}$ years or $\sim 45\times 10^{16}$ seconds. The time $T(n)$ we compute measures the number of steps of our integration algorithm. In order to compare with the age of the universe we have to convert it into seconds. For this we need the values of the parameters $\epsilon$ and $\sigma$ entering the definition of the Lennard Jones potential, the mass of the particles and the value of the discrete time steps $h$ we take in integrating the equations of motion. For every value of the density and the temperature, $h$ is kept constant, independent of the number of particles $N$. Assuming for $\epsilon$ and $\sigma$, the typical values of a rare gas interaction, for instance argon[7], we can estimate $N_{u}$. $N_{u}$ decreases when the gas density increases. We find that for $d=0.064$ $N_{u}\sim 294$ while for $d=0.191$ $N_{u}\sim 171$. Those numbers are very much smaller than $10^{23}$.
Thus our results overwhelmingly support Boltzmann’s point of view on the origin of irreversibility in statistical mechanics and there is no need for interactions violating time reversal symmetry.